in one year carla earned 460$ in simple interest when she invests 6200$. some of the money earned at an interest rate of 5% and the rest at 8% how much did she invest at each rate
Answer by mananth(16946)

Part I 8.00% per annum ------------- Amount invested =x
Part II 5.00% per annum ------------ Amount invested = y
6200
Interest----- 460.00

Part I 8.00% per annum ---x
Part II 5.00% per annum ---y
Total investment
x + 1 y= 6200 -------------1
Interest on both investments
8.00% x + 5.00% y= 460
Multiply by 100
8 x + 5 y= 46000.00 --------2
Multiply (1) by -8
we get
-8 x -8 y= -49600.00
Add this to (2)
0 x -3 y= -3600
divide by -3
y = 1200 - x)
Part I 8.00% $ 5000
Part II 5.00% $ 1200

CHECK
5000 --------- 8.00% ------- 400.00
1200 ------------- 5.00% ------- 60.00
Total -------------------- 460.00

m.ananth@hotmail.ca